OpenStudy (anonymous):

Find the polynomial f(x) that has the roots of –2, 3 of multiplicity 2. Explain how you would verify the zeros of f(x).

OpenStudy (anonymous):

please help?

OpenStudy (anonymous):

please?

OpenStudy (anonymous):

\[(x+2)^2(x-3)^2\]

OpenStudy (anonymous):

multiplicity 2 than root is bound

OpenStudy (anonymous):

\[as( x) \rightarrow \infty, y \rightarrow \infty\] \[as(x)\rightarrow-\infty, y \rightarrow \infty\]

OpenStudy (anonymous):

well if u mean only 3 has the multiplicity 2 then it would be (x+2)(x-3)^2. If both are multiplicity 2 then it would be (x+2)^2 X (x-3)^2. u verify the zeroes by taking each thing under the parenthesis to zero so x+2 =0 and x-3= 0 so isolate x and we got x=-2 and x=3

OpenStudy (anonymous):

maybe only 3 has mult of 2 x+2)(x-3)^2

OpenStudy (anonymous):

f(x)= (x+2)(x-3)^2 f(x)= (x+2)(x-3)^2=0 x+2=0, x=2 and (x-3)^2=0, x=3 and 3